Disturbed Corp. needs to raise $56.5 million to fund a new project. The company will sell shares at a price of $23.60 in a general cash offer and the company's underwriters will charge a spread of 7 percent. The direct flotation costs associated with the issue are $700,000 and the indirect costs are $435,000. How many shares need to be sold? Multiple Choice 2,265,167 shares 2,510,024 shares 2,254,673 shares 2,625,980 shares 2,394,068 shares
Answer #1

net cash received value per share after spread

=23.60*(1-7%)

=21.9480

money to be raised=56.5*1000000+700000+435000=57635000

How many shares need to be sold=57635000/21.9480

=2625980 shares

the above is answer..